### v1

# Autocovariance and Optimal Design for Random Walk Metropolis-Hastings Algorithm

Abstract
The Metropolis–Hastings algorithm has been extensively studied in the estimation and simulation literature, with most prior work focusing on convergence behavior and asymptotic theory. However, its covariance structure—an important statistical property for both theory and implementation—remains less understood. In this work, we provide new theoretical insights into the scalar case, focusing primarily on symmetric unimodal target distributions with symmetric random walk proposals, where we also establish an optimal proposal design. In addition, we derive some more general results beyond this setting. For the high-dimensional case, we relate the covariance matrix to the classical 0.23 average acceptance rate tuning criterion.
Keywords:
Autocovariance, autocorrelation, Markov chain Monte Carlo (MCMC)
I Introduction
Markov chain Monte Carlo (MCMC) provides a general framework for sampling from complex distributions by constructing a Markov chain with the desired stationary distribution [Spall2003]. This approach enables efficient approximation of expectations and integrals that are otherwise analytically intractable, and it has found widespread applications in statistics, engineering, and control [HillandSpall]. In particular, Monte Carlo particle filters can discretize the state space to enable efficient MAP sequence estimation via dynamic programming [Godsill2001]. MCMC has also been applied to Bayesian inference for stochastic kinetic models by sampling joint posteriors of parameters and latent data [Golightly2005], and to system identification for Bayesian parameter estimation in complex dynamical models beyond the reach of traditional techniques [Ninness2010]. In the control domain, particle MCMC has been applied to infer latent trajectories and nonparametric dynamics in Gaussian process state-space models [Frigola2013], MCMC has been used to generate model uncertainty sets in nonlinear dynamical systems via Hamiltonian Monte Carlo sampling of high-dimensional transfer operators [Srinivasan], and ghost sampling, an MCMC-based technique, has been introduced to efficiently simulate power disturbances conditional on rate-of-change-of-frequency violations in power systems [Moriarty]; moreover, MCMC has also been integrated with Bernstein approximation techniques to handle non-affine and dependent chance constrained optimization problems, providing an efficient approach to high-dimensional integration [Zhao].
The Metropolis–Hastings algorithm is one of the foundational methods in MCMC, with widespread applications in model estimation, physics, and system identification [HillandSpall, Metropolis1953]. The theoretical study of MH has been extensive, with a large body of work devoted to convergence behavior and asymptotic properties of estimators [Tierney1994, RobertsandRosenthal, BrooksandGareth1998]. An important line of research in MCMC has focused on the scaling and efficiency of Metropolis-Hastings algorithms in high-dimensional settings. For Gaussian random walk Metropolis (RWM) algorithms, the efficiency of the chain is maximized when the average acceptance rate is tuned to approximately $0.23$ [RobertsandGelman]. This “$0.23$ rule” has since become a widely used guideline for practitioners, providing a concrete and interpretable criterion for tuning proposal variances in high-dimensional problems.
While convergence and asymptotics have been well developed, comparatively little is known about the covariance structure, or equivalently the autocorrelation, of the MH chain itself. MCMC methods are closely connected with statistical estimation, providing a means to approximate otherwise intractable quantities. Within this framework, the covariance structure of the samples plays a central role in determining estimation accuracy and efficiency. Further knowledge of the covariances of the samples relates strategies for variance reduction, and potentially guides practical implementation of MCMC algorithms.
This work aims to address this gap. We first study the scalar case of MH with symmetric random walk proposals, focusing on symmetric unimodal target distributions. In this setting, we establish new theoretical results on the covariance structure and propose an optimal design for the proposal distribution of symmetric random walk form. In addition, we present some general results beyond this restrictive setting. For the high-dimensional case, under the same setup as Roberts and Gelman [RobertsandGelman], we relate the unit-lag covariance matrix to the classical 0.23 average acceptance rate criterion, providing new theoretical connections between covariance properties and established tuning principles.
Before beginning the technical analysis, we emphasize that Metropolis-Hastings chains are not necessarily positively correlated. As a counterexample, consider the target density $\pi(x)=\mathcal{N}(0,1)$ and the proposal density $q(y|x)=\mathcal{N}(-cx,2)$ with $c\in(0,1)$. In this setting, proposed moves tend to push the chain in the opposite direction of the current state, so that $x_{t}>0$ typically leads to $x_{t+1}<0$ and vice versa. Since the covariance reduces to $\mathrm{Cov}(X_{t},X_{t+1})=\mathbb{E}[X_{t}X_{t+1}]$ under the symmetry of $\pi$, the systematic flipping effect yields negative correlation. Both theoretical analysis and numerical evaluations confirm that the covariance is strictly negative for all $c\in(0,1)$.
The paper is organized as follows. Section II presents the preliminary background, notation, and problem setup. Section III contains the main theoretical results. Section IV discusses the optimal proposal design. Section V provides numerical experiments to assess whether the proposed design can improve the efficiency of the Metropolis-Hastings algorithm. Section VI concludes with remarks and directions for future work.
II Preliminaries
Monte Carlo methods provide a general tool for estimation and simulation when analytical solutions are unavailable. MCMC extends this idea by constructing a Markov chain with the desired stationary distribution, enabling sampling from complex or high-dimensional models [Spall2003, HillandSpall]. Within this framework, the Metropolis-Hastings (MH) algorithm stands out as a versatile and widely used scheme. In this section, we present the Metropolis-Hastings algorithm, establish the notation, and outline the problem setup.
II-A Metropolis-Hastings algorithm
Assume a continuous state space with probability density functions. The overall aim of MCMC methods such as Metropolis-Hastings algorithm is to generate samples from a target density of interest, say $\pi(x)$. The MH algorithm is defined by a proposal density $q(y|x)$ for transitioning from state $x$ to state $y$, and a target density $\pi$ of interest. The MH algorithm accepts a proposed move with probability
$\alpha(x,y)=\min\!\left(1,\frac{\pi(y)q(x|y)}{\pi(x)q(y|x)}\right),$
and otherwise the chain remains at $x$. This induces a Markov transition kernel
$\mathcal{K}(x,y)=\alpha(x,y)q(y|x)+\delta(y-x)\big(1-r(x)\big),$
where $\delta(\cdot)$ is the Dirac measure at 0, and $r(x)=\int\alpha(x,y)\,q(y|x)\,\mathrm{d}y$ is the overall acceptance probability at current state [HillandSpall]. Under ergodicity and irreducibility assumptions, the Markov chain converges to the unique invariant distribution. Let $(\mathcal{X},\mathcal{B})$ denote the underlying measurable state space, with $\mathcal{X}$ the state space and $\mathcal{B}$ its associated $\sigma$-algebra, then
$\pi(y)=\int_{\mathcal{X}}\mathcal{K}(x,y)\,\pi(dx)\quad\pi\text{-a.s. on }(\mathcal{X},\mathcal{B}).$
where $\pi$-a.s.” indicates that the equality holds for all $y$ except on a $\pi$-null set (i.e. almost surely).
The above formulation specifies the acceptance mechanism and transition kernel in mathematical terms. For clarity, we now summarize the MH algorithm in a step-by-step procedure.
  1. Step 1.
    Initialization. Choose the initial state $X_{0}\in\mathcal{X}$ and set $t=0$.
  2. Step 2.
    Given $X_{t}$, generate a candidate $Y$ according to the proposal density $q(\cdot|X_{t})$.
  3. Step 3.
    Compute the acceptance probability
    $\alpha(X_{t},Y)=\min\!\left(1,\frac{\pi(Y)q(X_{t}|Y)}{\pi(X_{t})q(Y|X_{t})}\right).$
    Then set
    $X_{t+1}=\begin{cases}Y,&\text{with probability }\alpha(X_{t},Y),\\[6.0pt] X_{t},&\text{with probability }1-\alpha(X_{t},Y).\end{cases}$
    Increment $t\leftarrow t+1$.
  4. Step 4.
    Repeat Steps 2-3 until the terminal iteration $n$ is reached, yielding the chain $(X_{0},\dots,X_{n})$.
Our aim in this paper is to analyze the autocovariance structure of the sequence $\{X_{0},X_{1},X_{2},\ldots,X_{t}\}$ produced by the MH algorithm. We focus on the larget $t$ (post-“burn-in”) where the process has achieved a stationary probability distribution.
II-B Notation
Let $X_{t}$ denote the current state. We consider the continuous state spaces, where $X_{t+k}\in\mathbb{R}$. The target distribution $\pi(\cdot)$ is assumed to have mean $\mu$ and strictly positive finite variance $\sigma_{\pi}^{2}\in(0,\infty)$. The proposal distribution follows a random walk of the form $X_{t+1}=X_{t}+Z,\,Z\sim\phi(\cdot)$, where $\phi(\cdot)$ denotes a density symmetric about 0, not restricted to the Gaussian case.
Throughout this paper we assume that the target distribution $\pi$ admits a density with respect to Lebesgue measure. Similarly, the proposal distribution $q(\cdot|x)$ is assumed to admit a density $q(y|x)$ with respect to the corresponding base measure. $\Pi(\cdot)$ is the cumulative distribution function corresponding to $\pi(\cdot)$, and we will use the cdf notation for brevity in what follows. Unless otherwise stated, all integrals are taken over $\mathbb{R}$ with respect to Lebesgue measure. We also adopt the standard stochastic order notation: $Z_{n}=o_{p}(a_{n})$ means $Z_{n}/a_{n}\to 0$ in probability, while $Z_{n}=O_{p}(a_{n})$ means $Z_{n}/a_{n}$ is bounded in probability. Also for simplicity, let $\partial_{x}f(x)$ and $\partial_{x}^{2}f(x)$ be the first and second derivatives of a function $f(x)$ with respect to its argument. Lastly, for a matrix $A$, $\|A\|_{\mathrm{op}}\mathrel{\mathop{\ordinarycolon}}=\max\{\,\|Ax\|_{2}\mathrel{\mathop{\ordinarycolon}}x\in\mathbb{R}^{n},\ \|x\|_{2}=1\,\}$ denotes its operator (spectral) norm.
II-C Problem Setup
Consider the unit-lag covariance of the output sequence of the MH algorithm under the stationarity assumption, i.e., the current state $X_{t}\sim\pi(\cdot)$, which implies that the next state $X_{t+1}\sim\pi(\cdot)$ as well. Stationarity holds provided that the chain is irreducible, aperiodic, and satisfies detailed balance with respect to $\pi(\cdot)$, ensuring that $\pi(\cdot)$ is the unique invariant distribution [HillandSpall]; these conditions are known to hold for an MH algorithm.
Our goal is to investigate the multi-lag covariance of a symmetric random walk MH chain, with particular interest in the structure of the covariance and the optimal design of the proposal distribution. We begin with the unit-lag covariance, $\mathrm{Cov}(X_{t},X_{t+1})$.
III Main results
In this section, we present our main theoretical results, deriving an analytic expression for the unit-lag covariance of a Metropolis-Hastings chain with a symmetric random-walk proposal and a symmetric unimodal target. This serves as the foundation for the optimal proposal design studied in Section IV.
Lemma 3.1.
For a Metropolis-Hastings chain, the general formula for the unit-lag covariance is given by
$\mathrm{Cov}(X_{t},X_{t+1})$
$=\sigma_{\pi}^{2}-\tfrac{1}{2}\iint(x-y)^{2}\pi(x)q(y|x)\alpha(x,y)\,\mathrm{d}x\,\mathrm{d}y.$ (1)
Recall that, unless otherwise specified, all integrals are taken over $\mathbb{R}$.
Proof.
Using tower property,
$\mathrm{Cov}(X_{t},X_{t+1})$ $=\mathbb{E}[X_{t}X_{t+1}]-\mathbb{E}[X_{t}]\mathbb{E}[X_{t+1}]$
$=\mathbb{E}[\mathbb{E}[X_{t+1}|X_{t}]\cdot X_{t}]-\mathbb{E}[X_{t}]\mathbb{E}[X_{t+1}].$
By stationariy, we know $\mathbb{E}[X_{t}]=\mathbb{E}[X_{t+1}]=\mu$. Then computing $\mathbb{E}[X_{t+1}|X_{t}=x]$,
$\mathbb{E}[X_{t+1}|X_{t}=x]$ $=\int y\cdot\mathcal{K}(x,y)\,\mathrm{d}y$
$=\int y\cdot\alpha(x,y)q(y|x)\,\mathrm{d}y+x\cdot(1-r(x))$
$=x+\int(y-x)\alpha(x,y)q(y|x)\,\mathrm{d}y.$
Then by reversibility, i.e. $\pi(x)q(y|x)=\pi(y)q(x|y)$,
$\mathbb{E}[\mathbb{E}[X_{t+1}|X_{t}]\cdot X_{t}]$
$=\iint x(y-x)\alpha(x,y)q(y|x)\pi(x)\,\mathrm{d}y\,\mathrm{d}x+\int x^{2}\pi(x)\,\mathrm{d}x$
$=-\frac{1}{2}\iint(x-y)^{2}\alpha(x,y)q(y|x)\pi(x)\,\mathrm{d}y\,\mathrm{d}x+\int x^{2}\pi(x)\,\mathrm{d}x.$ (2)
Applying the Tonelli’s theorem [Tonelli] to justify the interchange of integrals in (III) since the integrand $(x-y)^{2}\alpha(x,y)q(y|x)\pi(x)$ is nonnegative, and using stationarity, we obtain (3.1). ∎
If we further suppose the proposal density is of random walk form (not necessarily symmetric), (3.1) simplifies to
$\mathrm{Cov}(X_{t},X_{t+1})$
$=\sigma_{\pi}^{2}-\frac{1}{2}\iint(x-y)^{2}\phi(y-x)\pi(x)\alpha(x,y)\,\mathrm{d}x\,\mathrm{d}y$ (3)
since the proposal density $q(\cdot|x)$ is of random-walk form, i.e., $Y=X+Z$ with $Z\sim\phi(\cdot)$, we have $q(y|x)=\phi(y-x)$ and $q(x|y)=\phi(x-y)$. Substituting these expressions into (3.1) yields the desired result.
Moreover, when the proposal density is of symmetric random-walk form (not necessarily Gaussian), that is, $q(y|x)=q(x|y)=\phi(|x-y|)$, equation (3.1) takes the following form
$\mathrm{Cov}(X_{t},X_{t+1})$
$=\sigma_{\pi}^{2}-\tfrac{1}{2}\iint(x-y)^{2}\phi(y-x)\min\{\pi(x),\pi(y)\}\,\mathrm{d}x\,\mathrm{d}y.$ (4)
Lemma 3.1 expresses the unit-lag covariance in the quadratic form $(x-y)^{2}$, which naturally connects to random-walk proposals and enables explicit analysis. Building on this, Corollary 3.2 follows when the proposal is further assumed to be of random-walk form.
Corollary 3.2.
Suppose the proposal density is random-walk (not necessarily symmetric), then the unit-lag covariance is invariant under translation of the target density. Define $\pi_{c}(x)\mathrel{\mathop{\ordinarycolon}}=\pi(x-c),\,c\in\mathbb{R}$, then
$\mathrm{Cov}_{\pi_{c}}(X_{t},X_{t+1})=\mathrm{Cov}_{\pi}(X_{t},X_{t+1}).$
Proof.
Recalling that the proposal $q(\cdot|x)$ is of the random walk form, $q(y|x)=\phi(y-x)$ and $q(x|y)=\phi(x-y)$. Thus by definition, the acceptance rate under the shifted target density $\pi_{c}(\cdot)$ is
$\alpha_{c}(x,y)$ $=\min\bigg\{1,\frac{\pi(y-c)\,\phi(x-y)}{\pi(x-c)\,\phi(y-x)}\bigg\}.$
Recall from (3.1) that $\mathrm{Cov}_{\pi_{c}}(X_{t},X_{t+1})$ is
$\mathrm{Cov}_{\pi_{c}}(X_{t},X_{t+1})$
$=\sigma_{\pi}^{2}-\tfrac{1}{2}\iint(x-y)^{2}\phi(x-y)\pi_{c}(x)\alpha_{c}(x,y)\,\mathrm{d}x\,\mathrm{d}y$
$=\sigma_{\pi}^{2}-\tfrac{1}{2}\iint(x-y)^{2}\phi(x-y)\pi(x-c)\alpha(x-c,y-c)\,\mathrm{d}x\,\mathrm{d}y$
$=\sigma_{\pi}^{2}-\tfrac{1}{2}\iint(s-t)^{2}\phi(s-t)\pi(s)\alpha(s,t)\,\mathrm{d}s\,\mathrm{d}t$
$=\mathrm{Cov}_{\pi}(X_{t},X_{t+1})$
where we applied the change of variables xc=sx-c=s, yc=ty-c=t, noting that the variance is invariant under translation. Hence we may simply denote $\mathrm{Cov}_{\pi_{c}}(X_{t},X_{t+1})$ by $\mathrm{Cov}_{\pi}(X_{t},X_{t+1})$. ∎
The argument in Corollary 3.2 extends naturally to higher lags. That is, for lag $k$, the $k$-step transition kernel remains invariant under translations of the target density $\pi$, so the covariance is also invariant under translations of the target density. Corollary 3.2 also allows us, without loss of generality, to assume a zero mean ($\mu=0$) when considering symmetric unimodal target densities later.
Lemma 3.3.
Let $\phi(x)$ be any probability density that is symmetric about 0. Then,
$|x|\phi(x)<\frac{1}{2}\text{ a.s.}$ (5)
Proof.
Assume $\exists\,x_{0}\in\mathbb{R}$ s.t.
$|x_{0}|\phi(x_{0})\geq\frac{1}{2}\Rightarrow\phi(x_{0})\geq\frac{1}{2|x_{0}|}.$
By symmetry, $\phi(-x_{0})=\phi(x_{0})$, so
$\int_{-|x_{0}|}^{|x_{0}|}\phi(x)\,dx\geq\int_{-|x_{0}|}^{|x_{0}|}\frac{1}{2|x_{0}|}\,dx=1.$
Since $\phi$ is a probability density, equality must hold and $\phi(x)=\frac{1}{2|x_{0}|}$ a.s. on $|x_{0}|,|x_{0}|]$, i.e., $\phi$ is uniform, that is,
$|x|\phi(x)=\frac{|x|}{2|x_{0}|}<\frac{1}{2},\quad\forall\,|x|<|x_{0}|.$
At $x=\pm|x_{0}|$, equality may hold pointwise, but probability densities are defined up to sets of measure zero. We reached a contradiction. ∎
Lemma 3.3 is powerful in that it holds for any symmetric probability density, including standard distributions such as the normal $\mathcal{N}(0,1)$, for which $\max_{x}\{|x|\phi(x)\}=e^{-\frac{1}{2}}/{\sqrt{2\pi}}<1/2$. Its general validity makes it an essential tool for the proof of our main theorem.
Theorem 3.4.
Consider a Metropolis-Hastings chain with a symmetric random-walk proposal density and a symmetric unimodal target density with mean $\mu$. Let the centered cumulative distribution function be defined by $\Pi_{\mu}(x)\mathrel{\mathop{\ordinarycolon}}=\Pi(x+\mu)$. Then the unit-lag covariance admits the explicit representation
$\mathrm{Cov}(X_{t},X_{t+1})=4\int_{0}^{\infty}x\,[1-\Pi_{\mu}(x)]\,[1-4x\phi(2x)]\,\mathrm{d}x.$ (6)
Proof.
First define
$f(x)\mathrel{\mathop{\ordinarycolon}}=\int\min\{\pi(x+y),\pi(y)\}\,\mathrm{d}y.$
By Corollary 3.2, we may assume without loss of generality that $\mu=0$, then consider f(x)f(-x)
$f(-x)$ $=\int\min\{\pi(-x+y),\pi(y)\}\,\mathrm{d}y$
$\overset{y\mapsto-y}{=}\int\min\{\pi(x+y),\pi(-y)\}\,\mathrm{d}y=f(x)$
Hence $f$ is even, i.e., f(x)=f(x)f(x)=f(-x), $\forall x\in\mathbb{R}$.
Recall from equation (III) that
$\mathrm{Cov}(X_{t},X_{t+1})$
$=\sigma_{\pi}^{2}-\tfrac{1}{2}\iint(x-y)^{2}\phi(y-x)\min\{\pi(x),\pi(y)\}\,\mathrm{d}x\,\mathrm{d}y$
$=\sigma_{\pi}^{2}-\tfrac{1}{2}\int x^{2}\phi(x)\int\min\{\pi(x+y),\pi(y)\}\,\mathrm{d}y\,\mathrm{d}x$
$=\sigma_{\pi}^{2}-\int_{0}^{\infty}x^{2}\phi(x)\int\min\{\pi(x+y),\pi(y)\}\,\mathrm{d}y\,\mathrm{d}x.$ (7)
where the second equality follows from a change of variables together with the Tonelli’s theorem [Tonelli] and the third equality follows from the fact that $f(\cdot)$ is even.
Given the target density is symmetric unimodal, now consider $f(x),\,x\geq 0$
$f(x)$ $=\int_{-\infty}^{-x/2}\pi(y)\,\mathrm{d}y\;+\;\int_{-x/2}^{\infty}\pi(x+y)\,\mathrm{d}y,\qquad x\geq 0$
$=\Pi(-x/2)+1-\Pi(x/2)$
$=2[1-\Pi(x/2)].$
Hence,
$\mathrm{Cov}(X_{t},X_{t+1})$ $=\sigma_{\pi}^{2}-2\int_{0}^{\infty}x^{2}\phi(x)[1-\Pi(x/2)]\,\mathrm{d}x$
$=\sigma_{\pi}^{2}-16\int_{0}^{\infty}x^{2}\phi(2x)[1-\Pi(x)]\,\mathrm{d}x.$ (8)
We can also express the target variance $\sigma_{\pi}^{2}$ via integration by parts as
$\sigma_{\pi}^{2}=2\int_{0}^{\infty}x^{2}\pi(x)\,\mathrm{d}x=4\int_{0}^{\infty}x\,[1-\Pi(x)]\,\mathrm{d}x.$ (9)
Substituting (9) into (III) yields (6).
Now we can readily extend the result to the case of any arbitrary symmetric unimodal density with nonzero mean.
By Corollary 3.2, proposal densities of symmetric random–walk form $q(y|x)=\phi(|y-x|)$ is transition-invariant. Then the covariance admits the same form as in the mean–zero case, with $\Pi$ replaced by $\Pi_{\mu}$. ∎
Having established the necessary preliminaries, we are now ready to present the main theorem.
Theorem 3.5.
For a Metropolis-Hastings chain with a symmetric random-walk proposal density and a symmetric unimodal target density, the unit-lag covariance is strictly positive.
Proof.
Combining Lemma 3.3 and Theorem 3.4 yields
$\mathrm{Cov}(X_{t},X_{t+1})=4\int_{0}^{\infty}x\,[1-\Pi_{\mu}(x)]\,[1-4x\phi(2x)]\,\mathrm{d}x>0$
This conclusion holds under our standing assumption that the target distribution is non-degenerate, i.e. $\sigma_{\pi}^{2}>0$. ∎
We emphasize that MH chains are not necessarily positively correlated. If the proposal density deviates from the symmetric random-walk form, then even under a symmetric unimodal target distribution, the resulting chain can exhibit strictly negative unit-lag correlation, as illustrated by the counterexample mentioned in Section I.
IV Optimal Proposal Design
Based on Theorem 3.4, we can carry out an optimal proposal design (i.e.  choose the best density $q(y|x)$). We define the optimal proposal design as the choice of a symmetric density $\phi$ associated with the proposal $q(y|x)=\phi(y-x)$ that minimizes the unit-lag covariance of the Metropolis-Hastings chain under a symmetric unimodal target. This criterion is motivated by the fact that the unit-lag covariance is the leading term in the asymptotic variance expansion of ergodic averages, thereby directly governing estimation efficiency. Specifically, recall (6)
$\mathrm{Cov}(X_{t},X_{t+1})$
$=4\int_{0}^{\infty}x\,[1-\Pi_{\mu}(x)]\,[1-4x\phi(2x)]\,\mathrm{d}x$
$=4\int_{0}^{\infty}x\,[1-\Pi_{\mu}(x)]\,\mathrm{d}x-16\int_{0}^{\infty}x^{2}\phi(2x)[1-\Pi_{\mu}(x)]\,\mathrm{d}x.$ (10)
Only the second term depends on $\phi(\cdot)$, and this dependence is linear. Thus, minimizing $\mathrm{Cov}(X_{t},X_{t+1})$ is equivalent to maximizing the second term. If we do not impose the standard regularity condition—namely that the support of the target is contained in the support of the proposal, $\mathrm{Supp}[\pi(\cdot)]\subseteq\mathrm{Supp}[q(\cdot|x)]$, which ensures the chain is irreducible — then the optimal $\phi(\cdot)$ takes the form of a two-point measure, as occurs when the first-order condition admits a unique solution (true for log-concave symmetric unimodal densities)
$\phi^{*}(x)=\tfrac{1}{2}\,\delta(x-x^{*})+\tfrac{1}{2}\,\delta(x+x^{*}),$
where $\delta(\cdot)$ is the Dirac measure at 0, $x^{*}=2y^{*}$ and
$y^{*}\in\arg\max_{y\geq 0}\,y^{2}[1-\Pi_{\mu}(y)].$
Note that $\arg\max$ may define a set, though for symmetric unimodal target densities it commonly reduces to a unique maximizer.
Now we check if such $y^{*}$ exists. First, the finite variance assumption implies $\mathbb{E}[X^{2}]<\infty$, which ensures
$\lim_{y\to\infty}w(y)=\lim_{y\to\infty}y^{2}\bigl(1-\Pi(y)\bigr)=0.$ (11)
Indeed, since $1-\Pi(x)=\mathbb{P}(X>x)\leq\mathbb{E}[X^{2}]/x^{2}$ by Markov’s inequality, the limit can only be 0, a strictly positive constant, or $\infty$. The latter two would force $\mathbb{E}[X^{2}]=\infty$, contradicting the finite variance assumption.
From (11), and noting that $w(0)=0$, $w(y)>0$ for all $y>0$, $\lim_{y\to\infty}w(y)=0$, and $w(\cdot)$ is continuous, the Extreme Value Theorem guarantees that both the maximum and the minimum (zero) exist. Hence $w_{\max}\mathrel{\mathop{\ordinarycolon}}=w(y^{*})$ must exist for any non-degenerate target distribution with finite variance.
Equivalently, $y^{*}$ can be determined by the first-order condition
$2[1-\Pi_{\mu}(y)]=y\pi_{\mu}(y),$ (12)
with the additional requirement that the second derivative at $y^{*}$ is negative to ensure optimality.
Remark 4.1.
equation (12) has a unique solution when the target density $\pi(\cdot)$ is log-concave. In this case, define $g(y)\mathrel{\mathop{\ordinarycolon}}=\pi(y)/(1-\Pi(y))$, which is monotone increasing, implying that $h(y)=yg(y)$ is strictly increasing. Hence the equation $h(y)=2$ has a unique solution, and since $w(0)=w(\infty)=0$. Moreover, this critical point is necessarily the global maximizer of $w(y)$. Indeed, if it were a saddle point, then since $w(x)$ is continuously differentiable, additional critical points would have to exist, which contradicts the uniqueness.
However, under the regularity condition, the optimization problem becomes an infimum problem rather than a maximization problem. In this case, the maximum achieved by the two-point measure cannot be attained, since the regularity condition is violated. We can prove this rigorously, as shown in Proposition 4.2.
Proposition 4.2.
Recall the second term in equation (IV). Let $w(y)\mathrel{\mathop{\ordinarycolon}}=y^{2}[1-\Pi_{\mu}(y)]$ for $y\geq 0$ and define
$\mathcal{J}(\phi)\mathrel{\mathop{\ordinarycolon}}=4\int_{0}^{\infty}w(s)\,\phi(2s)\,\mathrm{d}s,$
for any symmetric proposal density $\phi$ about 0. Let $w_{\max}=\sup_{y\geq 0}w(y)$ and $\mathcal{M}=\{y\geq 0\mathrel{\mathop{\ordinarycolon}}\,w(y)=w_{\max}\}$, which is nonempty. If $\phi$ is allowed to be a symmetric probability measure, the supremum $\sup_{\phi}\mathcal{J}(\phi)=w_{\max}$ is attained by putting all mass at $\pm 2y$ with $y\in\mathcal{M}$. If $\phi$ is required to be a continuous density and $\mathcal{M}$ consists only of isolated points, then the supremum is not attained; instead, one can construct a sequence of increasingly concentrated densities around $\{\pm 2y\mathrel{\mathop{\ordinarycolon}}\,y\in\mathcal{M}\}$ that makes $\mathcal{J}(\phi)$ arbitrarily close to $w_{\max}$.
Proof.
By symmetry, with $h(y)\mathrel{\mathop{\ordinarycolon}}=2\phi(2y)$ we have
$\mathcal{J}(\phi)=2\int_{0}^{\infty}w(s)h(s)\,\mathrm{d}s\leq 2\sup_{s\geq 0}w(s)\!\int_{0}^{\infty}h(s)\,\mathrm{d}s=w_{\max}$
with equality iff $h$ is supported on $\mathcal{M}$. Thus any symmetric atomic measure supported on $\{\pm 2y\mathrel{\mathop{\ordinarycolon}}y\in\mathcal{M}\}$ achieves $w_{\max}$.
If $\phi$ is a probability density, then the induced function $h(y)=2\phi(2y)$ is continuous in the sense that it cannot concentrate all its probability mass at finitely many points. The set $\mathcal{M}$ consists only of isolated points. Then away from $\mathcal{M}$, the continuity of $w(\cdot)$ guarantees that there exists $\varepsilon>0$ such that $w_{\max}-w(y)\geq\varepsilon$ for all $y$ sufficiently far from $\mathcal{M}$. Since any density $h$ must assign positive probability to regions outside arbitrarily small neighborhoods of $\mathcal{M}$, the weighted average $\int_{0}^{\infty}w(s)h(s)\,\mathrm{d}s\leq w_{\max}/2-\varepsilon/2<w_{\max}/2$, and hence $\mathcal{J}(\phi)<w_{\max}$.
Nevertheless, the two-point measure can be approximated arbitrarily well by densities concentrated near the $\{\pm 2y\mathrel{\mathop{\ordinarycolon}}y\in\mathcal{M}\}$, such as narrow Gaussians. As the concentration increases, the corresponding values of $\mathcal{J}(\phi)$ approach $w_{\max}$. Hence $\sup\mathcal{J}=w_{\max}$, but the maximum is not attained within the class of densities when the maximizers are isolated points.
Thus, when $\phi(\cdot)$ is required to be a probability density (by the regularity condition), minimizing the unit-lag covariance becomes an infimum problem. Nevertheless, the value $w_{\max}\mathrel{\mathop{\ordinarycolon}}=\sup_{y\geq 0}w(y)$, where $w(y)=y^{2}[1-\Pi(y)]$ and $\Pi$ is the target cumulative distribution function, can be approximated arbitrarily well. For instance, if $x^{*}=2y^{*}$ with $y^{*}\in\arg\max_{y\geq 0}w(y)$, then letting the variance parameter $\sigma^{2}\to 0$ in the Gaussian-component bimodal density
$\phi(x)\;=\;\tfrac{1}{2}\,\mathcal{N}(x^{*},\sigma^{2})\;+\;\tfrac{1}{2}\,\mathcal{N}(-x^{*},\sigma^{2})\to\phi^{*}\text{ as }\sigma\to 0$
yields convergence to the two-point measure $\phi$.
Continuing with the analysis of the unit-lag covariance structure, we next relate it to the “0.23 rule” for random-walk Gaussian Metropolis algorithms via the unit-lag covariance matrix. This classic criterion arises from diffusion limit analysis in high dimensions [RobertsandGelman] and prescribes tuning the average acceptance probability to about $0.23$ for optimal efficiency. Here we reinterpret this criterion through the lens of the unit-lag covariance matrix. We now state Theorem 4.3, following the notation and the setup of Gelman and Roberts [RobertsandGelman].
Theorem 4.3.
Let $d\in\mathbb{N}$ and consider the product target
$\pi_{d}(\boldsymbol{x})=\prod_{i=1}^{d}p_{i}(x_{i}),\qquad\boldsymbol{x}=(x_{1},\dots,x_{d})^{\top}\in\mathbb{R}^{d},$
where each $p_{i}$ is a strictly positive $C^{2}$ density on $\mathbb{R}$ with $\sigma_{i}^{2}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Var}_{p_{i}}(X_{i})\in(0,\infty),\,m_{i}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{p_{i}}\big[(\partial_{x}\log p_{i}(X_{i}))^{2}\big]\in(0,\infty),$ and $\mathrm{E}_{p_{i}}[|\partial_{x}\log p_{i}(X_{i})|^{3}],\,\mathrm{E}_{p_{i}}[|\partial_{x}^{2}\log p_{i}(X_{i})|]<\infty$ for all $i$. Consider a stationary random-walk Metropolis chain with the symmetric random walk Gaussian proposal density, i.e. $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{z}$, where $\boldsymbol{z}\sim\mathcal{N}\!\Big(\boldsymbol{0},\tfrac{\ell^{2}}{d}\,\boldsymbol{I}_{d}\Big)$, $\forall\ell>0$, and $\boldsymbol{I}_{d}$ denotes the $d\times d$ identity matrix. Define the averages $\bar{\sigma}_{d}^{2}\mathrel{\mathop{\ordinarycolon}}=d^{-1}\sum\limits_{i=1}^{d}\sigma_{i}^{2},\,\bar{m}_{d}\mathrel{\mathop{\ordinarycolon}}=d^{-1}\sum\limits_{i=1}^{d}m_{i}.$ Denote by $\bar{m}\mathrel{\mathop{\ordinarycolon}}=\lim\limits_{d\to\infty}\bar{m}_{d}$, assuming the limit exists in $(0,\infty)$. Then the unit-lag covariance matrix satisfies
$\mathrm{Cov}(\boldsymbol{x}_{t},\boldsymbol{x}_{t+1})$
$=\Big(\mathrm{diag}(\sigma_{1}^{2},\dots,\sigma_{d}^{2})-\frac{\ell^{2}}{2d}\cdot 2\,\Phi\!\big(-\tfrac{\ell\sqrt{\bar{m}}}{2}\big)\,\boldsymbol{I}_{d}\Big)\;+\;\boldsymbol{R}_{d}$ (13)
where $\boldsymbol{R}_{d}$ is the remainder matrix whose diagonal entries are $o(d^{-1})$ and off-diagonal entries are $O(d^{-2})$, so that $\|\boldsymbol{R}_{d}\|_{\mathrm{op}}=o(d^{-1})$, where $\|\cdot\|_{\mathrm{op}}$ denotes the operator (spectral) norm, chosen since it directly controls eigenvalues and hence positive definiteness. In particular, the diagonal entries
$\mathrm{Cov}(\boldsymbol{x}_{t},\boldsymbol{x}_{t+1})_{ii}=\sigma_{i}^{2}\;-\;\frac{\ell^{2}}{2d}\,2\,\Phi\!\Big(-\tfrac{\ell\sqrt{\bar{m}}}{2}\Big)\;+o(d^{-1})$
are strictly positive for all sufficiently large d and the matrix $\mathrm{Cov}(\boldsymbol{x}_{t},\boldsymbol{x}_{t+1})$ is positive definite.
Proof.
By Gelman and Roberts and apply the change of variable, let $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{z}$,
$\lim_{d\to\infty}\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})]=2\,\Phi\!\big(-\tfrac{\ell\sqrt{\bar{m}}}{2}\big)$
and $\boldsymbol{z}\text{ is independent of }\boldsymbol{x}$ which implies, for finite large $d$
$\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})]=2\,\Phi\!\big(-\tfrac{\ell\sqrt{\bar{m}}}{2}\big)+o(1).$ (14)
Following the same idea as Lemma 3.1, we can write the unit-lag covariance matrix as
$\mathrm{Cov}(\boldsymbol{x}_{t},\boldsymbol{x}_{t+1})=\boldsymbol{\Sigma}_{\pi}-\tfrac{1}{2}\,\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})\,\boldsymbol{z}\boldsymbol{z}^{\top}]$ (15)
where $\boldsymbol{\Sigma}_{\pi}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{\pi}\!\big[(\boldsymbol{x}-\boldsymbol{\mu})(\boldsymbol{x}-\boldsymbol{\mu})^{\top}\big]$ is the covariance matrix of the stationary distribution with mean $\boldsymbol{\mu}=\mathbb{E}_{\pi}[\boldsymbol{x}]$.
Then fix $\boldsymbol{x}$ and set $f(\boldsymbol{x,z})\mathrel{\mathop{\ordinarycolon}}=\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})$. We look at $\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})\boldsymbol{z}\boldsymbol{z}^{\top}]$, for $\boldsymbol{z}\sim\mathcal{N}(\boldsymbol{0},\ell^{2}\boldsymbol{I}_{d}/d)$, Applying Stein’s Lemma [stein] and the assumption $\mathrm{E}_{p_{i}}[|\partial_{x}^{2}\log p_{i}(X_{i})|]<\infty$ gives,
$\mathbb{E}[Z_{i}Z_{j}f(\boldsymbol{x,z})|\boldsymbol{x}]=\frac{\ell^{4}}{d^{2}}\,\mathbb{E}[\partial_{ij}f(\boldsymbol{x,z})|\boldsymbol{x}],$
$\mathbb{E}[Z_{i}^{2}f(\boldsymbol{x,z})|\boldsymbol{x}]=\frac{\ell^{2}}{d}\,\mathbb{E}[f(\boldsymbol{x,z})|\boldsymbol{x}]+\frac{\ell^{4}}{d^{2}}\,\mathbb{E}[\partial_{ii}f(\boldsymbol{x,z})|\boldsymbol{x}].$
By the tower property, taking full expectation yields
$\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})Z_{i}Z_{j}]=O(d^{-2})$ (16)
$\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})Z_{i}^{2}]=\frac{\ell^{2}}{d}\,\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})]+O(d^{-2})$ (17)
Note that $\boldsymbol{\Sigma}_{\pi}=\mathrm{diag}(\sigma_{1}^{2},\dots,\sigma_{d}^{2})$. Using (16) and(17),
$\mathbb{E}\big[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})\,\boldsymbol{z}\boldsymbol{z}^{\top}\big]=\frac{\ell^{2}}{d}\,\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})]\,\boldsymbol{I}_{d}-2\boldsymbol{R}_{d},$ (18)
also note that
$O(d^{-2})+o(d^{-1})=o(d^{-1})$
so substituting (14) and (18) into (15) yields (4.3) with a remainder matrix $\|\boldsymbol{R}_{d}\|_{\mathrm{op}}=o(d^{-1})$. Now we can write the diagonal entries as
$\mathrm{Cov}(\boldsymbol{x}_{t},\boldsymbol{x}_{t+1})_{ii}=\sigma_{i}^{2}\;-\;\frac{\ell^{2}}{2d}\,2\,\Phi\!\Big(-\tfrac{\ell\sqrt{\bar{m}}}{2}\Big)\;+o(d^{-1})$ (19)
which is strictly positive for all large $d$ since $\sigma_{i}^{2}>0$.
Note that the leading term $\boldsymbol{\Sigma}_{\pi}-\ell^{2}/d\cdot\Phi\!\Big(-\ell\sqrt{\bar{m}}/2\Big)\boldsymbol{I}_{d}$ in (4.3) is a real diagonal matrix, hence Hermitian and the remainder $\boldsymbol{R}_{d}$ arises from
$\tfrac{1}{2}\Big(\,\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})\,\boldsymbol{z}\boldsymbol{z}^{\top}]-\tfrac{\ell^{2}}{d}\,\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})]\boldsymbol{I}_{d}\Big),$
it follows that $\boldsymbol{R}_{d}$ is real symmetric, hence Hermitian as well. Therefore both the main diagonal term and the perturbation matrix $\boldsymbol{R}_{d}$ are Hermitian. Then by Weyl’s inequality [Horn_Johnson_2012], for Hermitian matrices $\boldsymbol{A}$ and $\boldsymbol{E}$,
$\lambda_{\min}(\boldsymbol{A}+\boldsymbol{E})\;\geq\;\lambda_{\min}(\boldsymbol{A})-\|\boldsymbol{E}\|_{\mathrm{op}}.$
In our case $\boldsymbol{A}$ is the diagonal term and $\boldsymbol{E}=\boldsymbol{R}_{d}$, so
$\lambda_{\min}\!\big(\mathrm{Cov}(\boldsymbol{x}_{t},\boldsymbol{x}_{t+1})\big)\;\geq\;\min_{i}\!\Big\{\sigma_{i}^{2}-\tfrac{\ell^{2}}{d}\cdot\Phi(-\tfrac{1}{2}\ell\sqrt{\bar{m}})\Big\}-\|\boldsymbol{R}_{d}\|_{\mathrm{op}}.$
Since $\min\limits_{i}\sigma_{i}^{2}>0$, the subtracted terms are of order $O(d^{-1})$ and $o(d^{-1})$, hence strictly smaller than $\min\limits_{i}\sigma_{i}^{2}$ for sufficiently large $d$. This shows that the covariance matrix is positive definite in high dimension. ∎
Theorem 4.3 shows that in the high-dimensional regime, for a product-form target with non-identically distributed components, the diagonal entries of the unit-lag covariance admit the expansion as equation (19). Gelman and Roberts define the follwoing
[…]
Since $h(\ell)$ coincides with the efficiency criterion in the diffusion limit analysis of Gelman and Roberts, the optimizer $\ell^{*}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{\ell}h(\ell)=2.38/\sqrt{\bar{m}}$ yields the well-known optimal acceptance rate $\mathbb{E}[\alpha(x,y)]\approx 0.23$, see [RobertsandGelman], while simultaneously minimizing the diagonal entries of the unit-lag covariance matrix by Theorem 4.3.
V Numerical Study
Figure 1: ESS versus $\mathbb{E}[\alpha(x,y)]$ under matched wall-clock time (2s). Dashed gray shows standard Gaussian proposals; solid lines show bimodal proposals with $\sigma_{q}/x^{\star}=0.20,0.40$.
To test the efficiency of symmetric random walk proposal density associated with symmetric bimodal density $\phi(\cdot)$. We conduct numerical experiments with three symmetric, unimodal target families: (i) the Gaussian, (ii) the logistic, and (iii) the generalized hyperbolic secant (GHS). Their probability density functions are, respectively,
$\pi_{\mathrm{g}}(x|\mu)$ $=\frac{1}{\sqrt{2\pi}}\exp\!\left(-\frac{(x-\mu)^{2}}{2}\right),$ (20)
$\pi_{\mathrm{ell}}(x|\mu)$ $=\frac{e^{-(x-\mu)}}{\bigl(1+e^{-(x-\mu)}\bigr)^{2}}=\tfrac{1}{4}\,\operatorname{sech}^{2}\!\Bigl(\tfrac{x-\mu}{2}\Bigr),$ (21)
$\pi_{\mathrm{ghs}}(x|\alpha,\mu,\sigma)$ $=\frac{c_{\alpha}}{\sigma}\,\operatorname{sech}^{\alpha}\!\Bigl(\frac{\pi\,(x-\mu)}{2\sigma}\Bigr),\,\alpha,\sigma>0,$ (22)
where $\operatorname{sech}(z)=1/\cosh(z)$ and the normalizing constant $c_{\alpha}$ depends only on $\alpha$ and admits the closed form
$c_{\alpha}=\frac{\sqrt{\pi}}{2}\,\frac{\Gamma\!\bigl(\tfrac{\alpha+1}{2}\bigr)}{\Gamma\!\bigl(\tfrac{\alpha}{2}\bigr)}=\left(\frac{2}{\pi}\int_{-\infty}^{\infty}\operatorname{sech}^{\alpha}(t)\,\mathrm{d}t\right)^{-1}.$
All three families are symmetric about $\mu$ and unimodal. For the Gaussian and logistic targets with unit scale (without loss of generality), the cumulative distribution functions are $\Phi(x-\mu)$ and $(1+e^{-(x-\mu)})^{-1}$, respectively, while the GHS cdf has no closed form and is evaluated numerically. Unless otherwise specified, we set the Gaussian variance to $1$, the logistic scale to $1$, and $\sigma=1,\,\alpha=1$ for the GHS family.
Figure 2: Unit-lag correlation decreases with narrower bimodal proposals, approaching the two-point limit—the unattainable minimum under any continuous proposal density.
For the proposal distribution, we consider two cases: a Gaussian–component symmetric bimodal density $\phi(x)=\tfrac{1}{2}\mathcal{N}(x^{*},\sigma_{q}^{2})+\tfrac{1}{2}\mathcal{N}(-x^{*},\sigma_{q}^{2})$ with $\sigma_{q}>0$, and a standard Gaussian symmetric unimodal density. To ensure a fair comparison, we evaluate each MH chain in terms of effective sample size (ESS) and unit-lag autocorrelation, matched at the same (or nearly the same) average acceptance rate and running time (In practice, we fixed the wall-clock budge). The ESS is defined as
$\mathrm{ESS}\mathrel{\mathop{\ordinarycolon}}=\frac{N}{1+2\sum_{k=1}^{\infty}\rho_{k}},$
where $N$ is the total number of samples and $\rho_{k}$ is the lag-$k$ autocorrelation [DefofESS]. ESS measures the number of independent samples that the correlated MCMC output is equivalent to, thereby quantifying the impact of autocorrelation on estimation efficiency. Across target densities, the bimodal design typically yields higher ESS and lower unit-lag autocorrelation than the standard Gaussian proposal.
In Figure 1, we report the effective sample size for three target densities (Gaussian, Logistic, and GHS) under matched average acceptance probability and a runtime of 2 seconds. For the bimodal proposal, two cases with $\sigma_{q}/x^{*}=0.20$ and $\sigma_{q}/x^{*}=0.40$ were tested. In both cases, the narrow bimodal Gaussian yields substantially higher ESS and thus greater sampling efficiency than the standard Gaussian, consistent with concentrating proposal mass near the optimal jumps at $\pm 2x^{*}$, where $x^{*}$ maximizes $w(x)$. This advantage, however, diminishes as $\mathrm{E}[\alpha(x,y)]$ approaches $50\%$, since in this regime the chain already accepts nearly half of the proposed moves, and the marginal benefit of concentrating proposal mass near the optimal jumps becomes negligible.
Figure 3: Histograms of MH samples versus target densities for Gaussian $(\mu=0,\sigma=1)$, Logistic $(\mu=3,\sigma=1)$, and GHS $(\alpha=1,\mu=-7,\sigma=1)$. All chains ran for $\approx 1$s with bimodal proposal $(\sigma_{q}/x^{\ast}=0.40)$.
In Figure 2, consider the GHS target density with $\mu=0$, variance $1$, and shape parameter $\alpha=1.5$, the theoretical unit-lag correlation (equal here to the unit-lag covariance) converges, as $1/\sigma_{q}$ increases, i.e.  $\sigma_{q}$ decreases, to the minimum value obtained by substituting $\phi^{*}$ into equation (6). The empirical unit-lag autocorrelation of the MH chain closely tracks the theoretical curve and fluctuates around this minimum, as seen in Figure 2. These oscillations reflect numerical variation, which is more pronounced for small $\sigma_{q}$ due to higher sample correlation and lower effective sample size.
In Figure 3, we present histograms of MH samples against the true target densities for three cases (Gaussian, Logistic, and GHS) under the $\sigma_{q}/x^{*}=0.40$. The close alignment between the empirical histograms and the target curves confirms that the MH algorithm using the narrow bimodal density $\phi$ accurately recovers the stationary distribution, i.e. the target distribution.
VI Concluding Remarks and Future Work
Understanding the autocovariance structure of the Metropolis–Hastings algorithm is essential for assessing its efficiency. In this paper, we presented new theoretical results for random-walk MH chains targeting symmetric unimodal densities and introduced an optimal proposal design based on the unit-lag covariance, ultimately relating the analysis to high-dimensional settings in an asymmetric sense. Future work will focus on relaxing the restrictive assumptions on target distributions and extending the framework to multi-lag covariance structures.
Acknowledgment
The author thanks Dr. James A. Fill for an earlier proof of Lemma 3.3 and for proving the equation (9), which plays a crucial role in establishing Theorem 3.5.